OpenStudy (zyberg):

By what integer should 540 be divided, so that remainder would make 75% of quotient.

OpenStudy (zyberg):

I had found a way to solve this by writing an equation 180/N = y + x/N where x - remainder, y - quotient and N - the divisor that we need. However, after simplifying it (taking x to be 0.75y and adding the fractions) it turns out that I would need to put quite a lot different values into y and partially "guess" the answer. Is there a solution that doesn't require any random putting of 20 or so numbers and checking for each case?

ganeshie8 (ganeshie8):

Let "a" be the required integer. You have to solve : ``` 540 = a*q + r 4r = 3q ```

ganeshie8 (ganeshie8):

Multiply first equation by 4 in order to eliminate r : ``` 540*4 = 4aq + 4r ```

ganeshie8 (ganeshie8):

``` 540*4 = 4aq + 3q ```

ganeshie8 (ganeshie8):

``` 540*4 = q*(4a + 3) ```

ganeshie8 (ganeshie8):

Here is the question : What do you notice about an integer of form `4a+3` ? Is it divisible by 4 ?

OpenStudy (zyberg):

Thank you, @ganeshie8 I thought about going that way, however, when I got to the 540*4 = q*(4a + 3) I didn't know what else to do.... 4a + 3 isn't divisible by 4. The remainder is 3 or -1. So, with that information we could notice that q is divisible by 4: 540 * 4 congruent to q(4a + 3) (mod 4) 0 congruent to 3q (mod 4), so q must be a multiple of 4. However, what to do now?

ganeshie8 (ganeshie8):

@Zyberg that's right, but more importantly notice that \(4a+3\) is "not" divisible by \(4\).

OpenStudy (zyberg):

@ganeshie8 hm, but what would that allow me to do?

ganeshie8 (ganeshie8):

``` 540*4 = q*(4a + 3) ```

ganeshie8 (ganeshie8):

factor the left hand side and kick out powers of 4

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=factor+540*4

ganeshie8 (ganeshie8):

Can we now say that \(4a+3\) must be a divisor of \(\large 3^3*5\) ?

OpenStudy (zyberg):

No, we can't. Or, we shouldn't be able to as far as I can see.

ganeshie8 (ganeshie8):

Why not ?

OpenStudy (zyberg):

Okay, then we can. But I don't see why we can.

ganeshie8 (ganeshie8):

Let me give you a simple example

ganeshie8 (ganeshie8):

If \( x*y = a*b\) and if you know that \(x\) and \(a\) are coprime, is below true ? \(a\) must be a divisor of \(y\).

ganeshie8 (ganeshie8):

example : \(6*35 = 7*30 \) Clearly \(7\) is a divisor of \(35\)

OpenStudy (zyberg):

Ohh, I understand it now. Made up a few questions and found answers to them ;)

ganeshie8 (ganeshie8):

Good ``` 4^2*(3^3*5)= q*(4a + 3) ``` Since 4^2 and 4a+3 are coprime, we can say that 4a+3 is a divisor of 3^3*5

OpenStudy (zyberg):

Thanks! So, what should we do now? Is there any more simplification left or should we try putting some values?

ganeshie8 (ganeshie8):

We can simplify further

ganeshie8 (ganeshie8):

use that same fact again : when a number of form 4a+3 is divided by 4, we must get a remainder 3

ganeshie8 (ganeshie8):

Look at 3^3*5

OpenStudy (zyberg):

Same thing. Remainder 3.

ganeshie8 (ganeshie8):

When 5 is divided by 4, the remainder is 1. When 3 is divided by 4, the remainder is 3. When 3^2 is divided by 4, the remainder is 1. When 3^3 is divided by 4, the remainder is 3.

ganeshie8 (ganeshie8):

So 4a+3 has to be 3 or 3*5 or 3^3*5 (because only these combinations leave the remainder 3)

ganeshie8 (ganeshie8):

3 cases, shouldn't be hard ?

OpenStudy (zyberg):

Yeah, not hard at all :) Thank you very much!

ganeshie8 (ganeshie8):

Np :)

ganeshie8 (ganeshie8):

Looks I have missed one case, we get 4 cases right ? ``` So 4a+3 has to be 3 or 3^3 or 3*5 or 3^3*5 (because only these combinations leave the remainder 3) ```

OpenStudy (zyberg):

Well, 3^3 isn't a, so it didn't do much damage to miss it. I found that only a = 33 (from 4a + 3 = 3^3 * 5) solves the problem. Thanks again :)

ganeshie8 (ganeshie8):

looks good..